# From free idempotent monoids to free multiplicatively idempotent rigs

### v1

@@ Introduction
From a mathematical perspective, the subject of this paper is the study of multiplicatively idempotent rigs (mirigs, Definition 1.2) and more specifically free mirigs. Perhaps surprisingly, these are finite (Lemma 3.6). Accordinly, the main problem that we address in this paper is ‘How many elements does the free mirig on $n$ generators have?’ The original discussions went as far as establishing the size of the free mirig on two generators ($284$ elements, Example 3.31) via computational methods which would be difficult to scale up. Here we get as far as computing the size of the free mirig on three generators, developing enough understanding of these structures to be able to effectively carry out the computation by hand. To summarize, the sequence computed thus far is:
@@
In Section 2, we summarise results from a paper of Green and Rees [GR52] regarding idempotent monoids (also called bands), whose structure is essential to understanding mirigs, and deriving a representation of elements of idempotent monoids by a certain kind of binary tree (Section 2.2). From this perspective we examine their subsemigroups, and more particularly their replete subsemigroups (Definition 2.20) which we later use as part of the data representing an element of a free mirig; in the process, we extract some further combinatorial objects (leftmost and rightmost paths, Definition 2.27) simplifying the presentation of replete subsemigroups.
@@ Acknowledgements
The motivating example of a rig is that of the natural numbers, where addition and multiplication take their usual interpretations, but the multiplication operation in the rigs we will consider today will behave rather differently to the multiplication of natural numbers.
@@
Before turning to mirigs, which are the subject of this article, we should mention that Vienney chose to call these ‘Boolean rigs’. It’s worth examining why he chose this name and why we depart from it.
@@
Amongst rings, those most relevant in classical logic and lattice theory are Boolean rings, which are those rings in which $x^{2}=x$ for every element $x$. The two-element ring $\mathbb{Z}/2\mathbb{Z}$ has these properties. Any Boolean ring is commutative ($xy=yx$ for all $x,y$), which is demonstrated by first observing that $-1=(-1)^{2}=1$, then that
@@
To generalize Boolean rings, several incompatible definitions of ‘Boolean semiring’ have been introduced over the past sixty years, as discussed by Guzmán in [Guz92]. Guzmán’s own proposed definition comes from considering the variety of rigs generated by the ring $\mathbb{Z}/2\mathbb{Z}$ and the rig $\mathbb{B}$, which in particular contains both Boolean rings and Boolean algebras viewed as rigs (so ignoring the negation operation, although this can be recovered when it exists).
@@
Beware that other authors have claimed the term ‘Boolean semiring’ for distinct notions. For instance, Guterman decided in [Gut09] to define a Boolean semiring as a subsemiring of a powerset algebra (a Boolean algebra), a definition inconsistent with the usual definition of Boolean rings given above…
@@
With these examples, it became clear that the link with logic that justified the (already increasingly tenuous) association of the previous varieties with Boole is broken. The derived meet operation fails to be commutative as soon as the multiplication operation does, while the join operation fails to be idempotent in free mirigs, even with zero generators (see Example 3.1, where $1\vee 1=1+1+1\neq 1$). In other words, there is no longer a canonical ordering on mirigs that might endow their elements with a logical interpretation. As such, we have chosen not to associate the adjective ‘Boolean’ with our objects of study in the present article.
@@
Before we move on we should make a final point about our choice of terminology. Graham Manuell pointed out in the Zulip discussion that the term idempotent semiring is already used to refer to a rig in which the addition is idempotent, so $x+x=x$ for all $x$. This includes the tropical semiring of natural numbers where addition is the $\max$ operation and multiplication is the usual addition of natural numbers, which can cause some confusion at first sight.
It is for these reasons that we ended up with the terminology ‘mirig’, which is close to a name suggested by John Baez. We shall examine the intersection of the class of additively idempotent rigs with the class of mirigs in Section 4.
@@ 2 Idempotent monoids
The free (not necessarily commutative) rig on a given set of generators $A$ can be constructed in two steps as the free commutative monoid on the free monoid on $A$. Correspondingly, we can construct the free mirig on $A$ as (a quotient of) the free commutative monoid on the free idempotent monoid on $A$. A thorough understanding of free idempotent monoids will be indispensible for studying free mirigs, so we spend this section establishing all of the relevant theory.
@@ 2.5 Replete subsemigroups: from branches to paths
All subsemigroups of $T_{2}$ are replete.
@@
From Example 2.19, we thus have $42$ replete subsemigroups of $T_{2}$. However, this result belies the general case.
@@
From here on, we denote by $[(a_{1},\dotsc,a_{k})\rangle$ the set of right branches with rightmost path $(a_{1},\dotsc,a_{k})$.
@@
The set of right branches $[(a_{1},\dotsc,a_{k})\rangle$ has cardinality
@@
A set $\operatorname{rb}(S)$ of right branches of height $k$ is replete if and only if it is a union of rightmost path equivalence classes and for $s,t\in S$ with $\operatorname{rmp}(s)=(a_{1},\dotsc,a_{k})$ and $\operatorname{rmp}(t)=(a^{\prime}_{1},\dotsc,a^{\prime}_{k})$ and each index $1\leq j\leq k$ we have $[(a_{1},\dotsc,a_{k})\ast_{r}(a^{\prime}_{j},\dotsc,a^{\prime}_{k})\rangle\subseteq\operatorname{rb}(S)$.
@@
BAFEDC BAFEDC BAFEDC
BAFEDC BAFEDC BAFEDC
BAFEDC BAFEDC BAFEDC
BAFEDC
Table 3: A diagrammatic representation of the $22$ (inhabited) replete right branch sets. In each hexagon, nodes A, B, C, D, E, F respectively represent the minimal replete branch sets $[(a,b,c)\rangle$, $[(a,c,b)\rangle$, $[(c,a,b)\rangle$, $[(c,b,a)\rangle$, $[(b,c,a)\rangle$, $[(b,a,c)\rangle$ and each connected shape represents the replete set of branches generated by the minimal sets indicated in red which it contains.
@@
These are the values appearing in the ‘$3$-branch sets’ row of Table 2.5.
To complete the count, recall that for each $A\in\bar{\operatorname{\alpha}}(S)$ the leftmost and rightmost paths present in $S_{A}$ are independent, so we need only count the right branches for each possible alphabet and square the results before multiplying by the permutation factor, as we did for Table 2. After doubling for the identity and including the trees of smaller height, we obtain the grand total.
@@ 3 Mirigs
With all of this theory in place, we can at last change gears and move onto mirigs.
@@ 3.4 Thickets
Following Lemmas 3.21 and 3.23, we call a lonely tree with coefficient $1$ in a maximal thicket $f$ a straggler and say $f$ is without stragglers if it has none, meaning $P(f_{A})\neq 1$ for all $A\subseteq[n]$. We can decompose any maximal $f\in\mathbb{N}_{2,2}[T_{n}]$ as $f^{\mathrm{str}}+f^{\mathrm{w/o}}$ where the first term is the sum of the stragglers of $f$ and the latter is the maximal summand of $f$ without stragglers.
@@ 4.1 Various characteristics
For $R^{(2,1)}_{2}$ and $R^{(2,1)}_{3}$, we proceed as we did for $R_{2}$ and $R_{3}$, respectively, constructing a table to compute the sum by alphabets. See Tables 4.3 and 4.3.
@@
where we move the coefficient of $3$ outside the sum to deduce that $|R^{(1,2)}_{n}|$ is always a multiple of $3$. The story is simple for $n=0$ and $1$, recovering $\mathbb{N}_{1,2}$ and $\mathbb{N}_{1,2}[T_{n}]$, respectively (in the latter case, no further identifications are made by imposing $x^{2}=x$). For $n=2$ and $3$, we can directly reuse the counts in Tables 3.31 and 3.31. In so doing, we obtain:
@@
Implicit in Examples 4.2, 4.3 and 4.4 is the fact that imposing the characteristic identity leaves the set of alphabets of a maximal thicket invariant, since equivalent thickets can only be produced from (or absorbed into) trees which are summands of that thicket. As such, the arguments of Section 3.4 and 3.6 can be reapplied to arrive at the given conclusions. For the final characteristic $(0,2)$ this is no longer the case, since for any forest $f$ and tree $t$, $f\simeq^{\prime}f+t+t$. However, we have already seen in Section 1 that in this case we are working with Boolean rings, which are commutative, so all trees on the same alphabet become equivalent. Since the parity function of a complementary triple is respected by the reduction to characteristic $(2,0)$, we arrive at the following simple presentation.

### v2

# affil0affil0affiliationtext: LIPN – UMR 7030 CNRS, Université Sorbonne Paris Nord

@@ Introduction
From a mathematical perspective, the subject of this paper is the study of multiplicatively idempotent rigs (mirigs, Definition 1.2) and more specifically free mirigs. Perhaps surprisingly, these are finite (Lemma 4.6). Accordingly, the main problem that we address in this paper is ‘How many elements does the free mirig on $n$ generators have?’ The original discussions went as far as establishing the size of the free mirig on two generators ($284$ elements, Example 4.31) via computational methods which would be difficult to scale up. Here we get as far as computing the size of the free mirig on three generators, developing enough understanding of these structures to be able to effectively carry out the computation by hand. To summarize, the sequence computed thus far is:
@@
In Section 2, we summarise results from a paper of Green and Rees [GR52] regarding idempotent monoids (also called bands), whose structure is essential to understanding mirigs, and deriving a representation of elements of idempotent monoids by a certain kind of binary tree (Section 2.2).
Using the tree presentation, we examine the subsemigroups of free idempotent monoids in Section 3. More particularly, we aim to understand their replete subsemigroups (Definition 3.6) because these will be used later as part of the data representing elements of free mirigs. In the process, we extract some further combinatorial objects (leftmost and rightmost paths, Definition 3.13) simplifying the presentation of replete subsemigroups.
@@ Acknowledgements
The motivating example of a rig is that of the natural numbers, where addition and multiplication take their usual interpretations. However, the multiplication operation in the rigs we will consider here will behave rather differently to the multiplication of natural numbers.
@@
Before turning to mirigs, which are the subject of this article, we should mention that Vienney used the term ‘Boolean rigs’, which is still the name associated with this structure on the nLab at time of writing [BV24]. It’s worth examining why he chose this name and why we depart from it.
@@
Amongst rings, those most relevant in classical logic and lattice theory are Boolean rings: rings in which $x^{2}=x$ for every element $x$, such as the two-element ring $\mathbb{Z}/2\mathbb{Z}$. Any Boolean ring is commutative ($xy=yx$ for all $x,y$), which is demonstrated by first observing that $-1=(-1)^{2}=1$, then that
@@
To generalize Boolean rings, several incompatible definitions of ‘Boolean semiring’ have been introduced in the past sixty years, as discussed by Guzmán in [Guz92]. Guzmán’s own proposed definition comes from considering the variety of rigs generated by the ring $\mathbb{Z}/2\mathbb{Z}$ and the rig $\mathbb{B}$, which in particular includes both Boolean rings and Boolean algebras viewed as rigs (ignoring negation and subtraction, respectively, although they can be recovered when they exists).
@@
Beware that other authors have claimed the term ‘Boolean semiring’ for distinct notions. For instance, Guterman decided in [Gut09] to define a Boolean semiring as a subsemiring of a powerset algebra (a Boolean algebra), a definition inconsistent with the usual definition of Boolean rings given above.
@@
With these examples, it became clear that the link with logic that justified the (already increasingly tenuous) association of the earlier structures with Boole is broken. The derived meet operation fails to be commutative as soon as the multiplication operation does, while the join operation fails to be idempotent in free mirigs, even with zero generators (see Example 4.1, where $1\vee 1=1+1+1\neq 1$). Thus there is no longer a canonical ordering on mirigs that might endow their elements with a logical interpretation, and as such we have chosen not to associate the adjective ‘Boolean’ with our objects of study.
@@
Before moving on we should make a final point about our choice of terminology. Graham Manuell pointed out in the Zulip discussion that the term idempotent semiring is already used to refer to a rig in which the addition is idempotent, so $x+x=x$ for all $x$. This includes the tropical semiring of natural numbers where addition is the $\max$ operation and multiplication is the usual addition of natural numbers, which can cause some confusion at first sight.
It is for this reason that we selected the terminology ‘mirig’, which is close to a name suggested by John Baez. We shall examine the intersection of the class of additively idempotent rigs with the class of mirigs in Section 5.
@@ 2 Idempotent monoids
The free (not necessarily commutative) rig on a given set of generators $A$ can be constructed in two steps as the free commutative monoid on the free monoid on $A$. Correspondingly, we can construct the free mirig on $A$ as a quotient of the free commutative monoid on the free idempotent monoid on $A$. A thorough understanding of free idempotent monoids will be indispensible for studying free mirigs, so we spend this section establishing all of the relevant theory.
@@ 2.3 Products of trees
3 Subsemigroups of free idempotent monoids
As we shall see in Section 4.4, we can present elements of the free mirig on $[n]$ as formal sums of trees in $T_{n}$: thickets, Definition 4.10. By ignoring how often each tree appears, a thicket determines a subset of $T_{n}$. Observing that $u+v\sim(u+v)^{2}\sim u+v+uv+vu$ in a mirig, we can anticipate that a maximal thicket representing a given element of the free mirig (supposing such exists) will determine a subset which is closed under products of trees: a subsemigroup of $T_{n}$. For this reason, we pre-emptively study subsemigroups in the present section, beginning with subsemigroups of a particularly simple form, the uniform subsemigroups, in Section 3.1. This will reveal that in identifying subsemigroups we can treat the left- and right-hand branches of the constituent trees almost independently, a fact which we exploit in the remainder of the article.
In fact, the subsemigroup determined by a maximal thicket must be closed under ‘internal multiplication’, in the sense that we have $x(u+v)y\sim xuy+xvy+xuvy+xvuy$, so that if both $xuy$ and $xvy$ are (independently) present, then $xuvy$ and $xvuy$ must be too. This leads us to the notion of replete subsemigroups in Section 3.2. The parenthetical ‘independently’ leads to some subtleties to be examined in later sections.
Since this paper is about counting structures, and since being able to count subsemigroups is a prerequisite for counting elements of the free mirig later, we enumerate the subsemigroups we find along the way. We count the uniform subsemigrops of $T_{n}$ in Theorem 3.2 and (with much greater effort!) count the replete subsemigroups of $T_{n}$ up to $n=3$ (see Lemma 3.10 and Proposition 3.20).
@@ 3.2 Replete subsemigroups: from branches to paths
All $42$ subsemigroups of $T_{2}$ from Example 3.5 are replete. We can thus record also that $T_{0}$ has $2$ subsemigroups and $T_{1}$ has $4$.
@@
As the reader may have expected, this result belies the general case.
@@
From here on, we denote by $[a_{1},\dotsc,a_{k}\rangle$ the set of right branches with rightmost path $(a_{1},\dotsc,a_{k})$.
@@
The set of right branches $[a_{1},\dotsc,a_{k}\rangle$ has cardinality
@@
A set $\operatorname{rb}(S)$ of right branches of height $k$ is replete if and only if it is a union of rightmost path equivalence classes and for $s,t\in S$ with $\operatorname{rmp}(s)=(a_{1},\dotsc,a_{k})$ and $\operatorname{rmp}(t)=(a^{\prime}_{1},\dotsc,a^{\prime}_{k})$ and each index $1\leq j\leq k$ the branch equivalence class of the right-most path $(a_{1},\dotsc,a_{k})\ast_{r}(a^{\prime}_{j},\dotsc,a^{\prime}_{k})$ is contained in $\operatorname{rb}(S)$.
@@
These are the values appearing in the ‘$3$-branch sets’ rows of Table 3.2.
To complete the count, recall from Theorem 3.2 that for each $A\in\bar{\operatorname{\alpha}}(S)$ the leftmost and rightmost paths present in $S_{A}$ can be chosen independently, so we need only count the right branches for each possible alphabet and square the results before multiplying by the permutation factor, as we did for Table 2. After doubling for the identity and including the trees of smaller height, we obtain the grand total.
@@
BAFEDC BAFEDC BAFEDC
BAFEDC BAFEDC BAFEDC
BAFEDC BAFEDC BAFEDC
BAFEDC
Table 3: A diagrammatic representation of the $22$ (inhabited) replete right branch sets. In each hexagon, nodes A, B, C, D, E, F respectively represent the minimal replete branch sets $[a,b,c\rangle$, $[a,c,b\rangle$, $[c,a,b\rangle$, $[c,b,a\rangle$, $[b,c,a\rangle$, $[b,a,c\rangle$ and each connected shape represents the replete set of branches generated by the minimal sets indicated in red which it contains.
@@ 4 Mirigs
With all of this theory in place, we can at last move on to mirigs.
@@ 4.4 Thickets
Following Lemmas 4.21 and 4.23, we call a lonely tree with coefficient $1$ in a maximal thicket $f$ a straggler and say $f$ is without stragglers if it has none, meaning $P(f_{A})\neq 1$ for all $A\subseteq[n]$. We can decompose any maximal $f\in\mathbb{N}_{2,2}[T_{n}]$ as $f^{\mathrm{str}}+f^{\mathrm{w/o}}$ where the first term is the sum of the stragglers of $f$ and the latter is the maximal summand of $f$ without stragglers.
@@ 5.1 Various characteristics
For $R^{(2,1)}_{2}$ and $R^{(2,1)}_{3}$, we proceed as we did for $R_{2}$ and $R_{3}$, respectively, constructing a table to compute the sum by alphabets. See Tables 5.3 and 5.3.
@@
where we move the coefficient of $3$ outside the sum to deduce that $|R^{(1,2)}_{n}|$ is always a multiple of $3$. The story is simple for $n=0$ and $1$, recovering $\mathbb{N}_{1,2}$ and $\mathbb{N}_{1,2}[T_{n}]$, respectively (in the latter case, no further identifications are made by imposing $x^{2}=x$). For $n=2$ and $3$, we can directly reuse the counts in Tables 4.31 and 4.31. In so doing, we obtain:
@@
Implicit in Examples 5.2, 5.3 and 5.4 is the fact that imposing the characteristic identity leaves the set of alphabets of a maximal thicket invariant, since equivalent thickets can only be produced from (or absorbed into) trees which are summands of that thicket. As such, the arguments of Section 4.4 and 4.6 can be reapplied to arrive at the given conclusions. For the final characteristic $(0,2)$ this is no longer the case, since for any forest $f$ and tree $t$, $f\simeq^{\prime}f+t+t$. However, we have already seen in Section 1 that in this case we are working with Boolean rings, which are commutative, so all trees on the same alphabet become equivalent. Since the parity function of a complementary triple is respected by the reduction to characteristic $(2,0)$, we arrive at the following simple presentation.